OpenStudy (anonymous):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero.

OpenStudy (astrophysics):

You must find the derivative twice

OpenStudy (anonymous):

would that be 1/x ?

OpenStudy (anonymous):

I mean t its 1/t

OpenStudy (astrophysics):

Yes, that is the acceleration but notice how it says when the velocity is first 0 so what is the velocity?

OpenStudy (anonymous):

im not sure

OpenStudy (astrophysics):

Velocity is the first derivative of displacement, acceleration is the second, so mathematically \[v= \frac{ ds }{ dt }, a = \frac{ d^2s }{ dt^2 }\] so you must have got the first derivative before you got the second, what was it?

OpenStudy (anonymous):

(ln)3x +1

OpenStudy (astrophysics):

Good, \[v = \ln(3t)+1\] now set v = 0 and solve for t, \[0 = \ln(3t)+1\]

OpenStudy (astrophysics):

I'll start you off, \[\ln(3t)=-1\] any idea how to isolate t?

OpenStudy (anonymous):

no

OpenStudy (astrophysics):

What if we exponentiate both sides

OpenStudy (anonymous):

how

OpenStudy (astrophysics):

\[\huge e^{\ln(3t)}=e^{-1}\]

OpenStudy (anonymous):

and then what?

OpenStudy (astrophysics):

The left side cancels out the logarithm so we're left with \[3t = e^{-1}\]

OpenStudy (astrophysics):

Finish it off

OpenStudy (anonymous):

t=(e^-1)/3

OpenStudy (astrophysics):

Yes, and use the rule \[x^{-n} = \frac{ 1 }{ x^n }\]

OpenStudy (anonymous):

so 1/e divided by 3

OpenStudy (astrophysics):

\[t = \frac{ 1 }{ 3e }\]

OpenStudy (anonymous):

would that be the answer

OpenStudy (astrophysics):

No, that is your t, now we must plug that in for a that will give us our answer for a at v = 0

OpenStudy (anonymous):

1/ 1/3e?

OpenStudy (astrophysics):

Yes, that simplifies to a(t)=3e

OpenStudy (astrophysics):

\[\huge \frac{ a }{ \frac{ c }{ d } } = \frac{ ad }{ c }\]